OpenStudy (anonymous):

Find an equation in slope-intercept form of a line that has a slope of 5 and passes through point A(4,-5),

OpenStudy (anonymous):

put in point slope form y-y1=m(x-x1) y+5=5(x-4) distribute y+5=5x-20 minus 5 y=5x-25

OpenStudy (anonymous):

Thanks!